OpenStudy (anonymous):

MY QUESTION GOT DELETED DELETED QUESTION CONTINUED HERE Please help Find angle T if angle T is 20 more than four times its supplement

OpenStudy (anonymous):

left off from t = 20 +4(supplement) but I don't know what supplement is supposed to mean

OpenStudy (msmr):

I would set it up as: t = 20 + 4(180-t)

OpenStudy (anonymous):

why 180-t?

OpenStudy (anonymous):

As I mentioned, if you add the measure of an angle, and the measure of its supplement, you get 180

OpenStudy (mathmate):

Supplement = (180-T) four times its supplement = 4(180-T) 20 more than four times its supplement = 20 + 4(180-T) So if "if angle T is 20 more than four times its supplement" T=20 + 4(180-T) as Msmr has posted.

OpenStudy (anonymous):

so if you want the measure of the supplement of T, you subtract t from 180

OpenStudy (anonymous):

Mathmate has the right equation now. You can solve it :-D

OpenStudy (anonymous):

is it 152? wait what I don't understand

OpenStudy (anonymous):

that's it i'm dropping out of high school good bye

OpenStudy (anonymous):

I'M WARNING YOu!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

lol dont drop out

OpenStudy (anonymous):

but no one understands how to do this ;(

OpenStudy (msmr):

a couple people have explained it...which part is confusing?

OpenStudy (anonymous):

i mean i don't understand how to do this

OpenStudy (anonymous):

i didn't mean no one i mean't i dont

OpenStudy (anonymous):

" if you add the measure of an angle, and the measure of its supplement, you get 180" i don't understand this

OpenStudy (mathmate):

T=20+4(180-T) T=20+720-4T 5T=740 T=740/5=148 degrees. Please.... do not drop out of high school!!!

OpenStudy (anonymous):

Yeah don't drop out. That was just the definition of a supplement of an angle.

OpenStudy (anonymous):

You just have to know it :-D

OpenStudy (mathmate):